OpenStudy (anonymous):

(x+2)(x-3)/(x+4)(x+2)*(x+4)(x-5)/(x-5)(x+6) please help me solve this all the way

OpenStudy (anonymous):

\[\LARGE \frac{(x+2)(x-3)}{(x+4)(x+2)}\cdot \frac{(x+4)(x-5)}{(x-5)(x+6)}=\] \[\LARGE \frac{\cancel{(x+2)}(x-3)}{(x+4)\cancel{(x+2)}}\cdot \frac{(x+4)\cancel{(x-5)}}{\cancel{(x-5)}(x+6)}=\] \[\LARGE {x-3\over x+4}\cdot {x+4\over x+6 }=...\]

OpenStudy (anonymous):

things witch are the same up and down cross them and get rid of them